Ch 11 Homework 1 Saved Help QS 11-14 Dividends on cumulative preferred stock LO C2 14.32 points Green Planet Corp. has (a) 6,Green Planet Corp. has (a) 6,400 shares of cumulative 10% preferred stock with a $2 par value and (b) 21,400 shares of common stock with a $0.01 par value. During its first two years of operation, Green Planet declared and paid the following total cash dividends. Compute the dividends paid each year to each of the two classes of stockholders: preferred and common.
Answer #1

Solution

Green Planet

Computation of dividends paid each year to each of the two classes of stockholders: preferred and common –

Dividends to

Cumulative Preferred

Common

Year 1

$650

0

Year 2

$1,910

$230

Computations:

Dividends to preferred stockholders:

The preferred stock is cumulative, which means that the unpaid dividends in a year are paid in the subsequent year.

Preferred stock = 6,400 x $2 = $12,800

Dividends = 10% x 12,800 = $1,280

Preferred dividend paid in first year = $650

Cumulative Dividend to be paid in second year = second year dividend + unpaid dividend of first year

= $1,280 + (1,280 – 650) = 1,280 + 630 = $1,910

Total cash dividends paid in Year 2 = $2,140

Preferred dividends paid in Year 2 = $1,910

Since dividends to preferred stockholders are paid first, the remaining dividends to common stockholders = 2,140 – 1,910 = $230